OpenStudy (anonymous):

simplify completely 3x+18 ________ 18

OpenStudy (anonymous):

3x+18/18

OpenStudy (anonymous):

= 3(x+6)/18 = (x+6)/6

OpenStudy (anonymous):

simp (4x+16)/16

OpenStudy (anonymous):

4(x+4)/16 =(x+4)/4